OpenStudy (anonymous):

Select the coordinates of two points that are on the line x = 5. A. (5,0) and (0,5) B. (5,0) and (5,-2) C. (5,0) and (5,2) D. (0,5) and (0,2)

OpenStudy (mvpriest6):

what do you think it is

OpenStudy (anonymous):

I'm not sure, these confuse me, could you help me understand it?

OpenStudy (anonymous):

A

OpenStudy (mvpriest6):

@felavin please dont give direct answers we are supposed to help

OpenStudy (anonymous):

THAT IS SIMPLE

OpenStudy (mvpriest6):

I know but he didnt get it

OpenStudy (anonymous):

Can you help with another one?

OpenStudy (mvpriest6):

sure

OpenStudy (anonymous):

And I pretty much got it before, but I wanted to make sure.

OpenStudy (mvpriest6):

okay

OpenStudy (anonymous):



OpenStudy (anonymous):

YOU MUST BE CHECK THE CORDINATE WHERE ON THE LINE (X=5 )

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). A. y=1 over 3x+7 B. y=-3x +3 C. y= 1 over 3x +3 D. y= -3x+7

OpenStudy (anonymous):

Thank you @felavin